A force of 10N acts on a body of mass 2kg for 0.5s. What is the increase in momentum?
1
Expert's answer
2021-03-10T17:20:16-0500

According to the second Newton's law, the increase in momentum is:


"\Delta P = F\Delta t"


where "F = 10N" is the average force that acts during the time "\Delta t = 0.5s". Thus, obtain:


"\Delta P = 10N\cdot 0.5s = 5 \space kg\cdot m/s"

Answer. "5 \space kg\cdot m/s".